IYF Corporation manufactures miscellaneous parts for building construction and maintenance. IYF uses a normal job costing system. The system applies manufacturing overhead on the basis of direct labor cost. For managerial purposes, over- or underapplied overhead is written off to Cost of Goods Sold monthly. IYF hires interns to work in its Plant Accounting department and, as a part of its interview process, asks candidates to take a short quiz.

You are given the following journal entries for June. (Assume that entries are made only once each month.)

Account Titles Debit Credit
Work-in-Process Inventory (Direct Labor) 12,000
Wages Payable 12,000
Direct Material Inventory 14,800
Accounts Payable 14,800
Finished Goods Inventory 35,200
Work-in-Process Inventory 35,200
Cost of Goods Solda 44,000
Finished Goods Inventory 44,000

aThis entry does not include any over- or underapplied overhead. Over- or underapplied overhead is written off to Cost of Goods Sold once for the month. For June, the amount written off was 6 percent of overhead applied for June.

The Work-in-Process ending account balance on June 30 was twice the beginning balance. The Direct Material Ending Inventory balance on June 30 was $6,940 less than the beginning balance. The Finished Goods ending balance on June 30 was $3,080.

The June income statement shows Cost of Goods Sold of $44,576.

Required:

a. What was the Finished Goods beginning inventory on June 1?

b. How much manufacturing overhead was applied for June?

c. Overhead is applied on the basis of direct labor costs. What was the manufacturing overhead rate for June?

d. How much manufacturing overhead was incurred for June?

e. What was the Work-in-Process beginning inventory balance?

f. What was the Work-in-Process ending inventory balance?
Answer #1
Calculation Answer
a. Finished goods inventory at beginning on June 1 (44,000+3080-35,200) $11,880
b. Manufacturing overhead applied COGG in income statement = 44,576
COGS from Finished goods A/c = 44,000
Hence, excess COGS = 576 due to underapplied OH.
Underapplied overhead = 6% of total overhead
Total overhead = 576/6% = $ 9600
$9,600
c. rate of manufacturing overhead Direct labor cost = 12,000
overhead applied = 9,600
rate of application = 9,600/12,000 = $ 0.8
$0.80
d. manufacturing overhead incurred Manufacturing overhead applied = 9600
Under applied overheads = 576
Hence, manufacturing overhead incurred = 10176
$10,176
e. Work in process beginning balance Refer note 1 $8,140
f. Work in process ending balance Refer note 1 $16,280
Note -1
Beginning WIP = X
Ending WIP = 2X
WIP - T aacount
Beg. Bal.             X Finished goods transfer     35,200
Wages             12,000 Ending balance                    2X
Overhead         9600
X+12,000 +9600 = 2X +35200
hence, X = 8140
Hence, 2X = 16280

*if you have any doubt please let me know in the comment section. Please give your valuable feedback.